OpenStudy (anonymous):

Simplify square root parenthesis 1 minus cosine theta parenthesis times parenthesis 1 plus cosine theta parenthesis. |sin Θ| ±cos Θ ±tan Θ square root cosine theta

OpenStudy (ipwnbunnies):

\[\sqrt{1 - \cos \theta}* \sqrt{1+\cos \theta}\]

OpenStudy (ipwnbunnies):

Is that right? When multiplying square root, you can but both factors under the square root, and expand it. \[\sqrt{(1+\cos \theta)(1 - \cos \theta)}\]

OpenStudy (anonymous):

yes that is the right question but I think the awnser is plus or minus cos theta but im not sure

OpenStudy (ipwnbunnies):

Just expand out what's under the square root. You'll see.

OpenStudy (anonymous):

wouldn't the awnser be the squareroot of cos and the squareroot of -cos

OpenStudy (ipwnbunnies):

Nooo. What did you get when you expanded out the factors. ;-;

OpenStudy (ipwnbunnies):

Nope. Try again. ;3 Look again at your last step in the FOIL method

OpenStudy (anonymous):

1-cosTheta+cosTheta-costhetasquared

OpenStudy (ipwnbunnies):

Yesh. Now simplify what's under the square root. c: You'll get. \[\sqrt{1 - \cos^2 \theta}\] Now, you'll have to use this trig identity: \[\sin^2 \theta + \cos^2 \theta = 1\] \[\sin^2 \theta = 1 - \cos^2 \theta\]

OpenStudy (anonymous):

so is the awnser A

OpenStudy (ipwnbunnies):

Let's not get ahead of ourselves now. ;-; Do the final steps.

OpenStudy (anonymous):

\[\sqrt{\sin ^{2}}\]

OpenStudy (ipwnbunnies):

.-. And...yeah lol. \[\sqrt{x^2} = |x|\] Therefore, \[\sqrt{\sin^2 \theta} = |\sin \theta|\]

OpenStudy (anonymous):

ok thanks